OpenStudy (myersra):

if f(x)= 3x-6/x-2, what is the average rate of change of f(x) over the interval [6,8]?

OpenStudy (ask_if_im_a_ghost):

Avg rate of change = (f(8)-f(6))/(8-6)

OpenStudy (mathmale):

3x-6 I suspect you mean f(x)=---------- x-2 Is this correct? If so, please use parentheses around each the numerator and the denominator, so that there is no ambiguity regarding your intent.

OpenStudy (myersra):

options are 0 1 3 or 3.5

OpenStudy (ask_if_im_a_ghost):

It is essentially the slope between the two ponts.

OpenStudy (ask_if_im_a_ghost):

No f(8)=(3*8-6)/(8-2)=3 f(6)=(3*6-6)/(6-2)=3 f(8)-f(6)=0 8-6=2 So the slope is 0/2=0 You could also realise that the function f(x) simplifies to 3, f(x)=3 has a slope of 0

OpenStudy (myersra):

thank you so much